What is the largest value of X makes this number 577,5X7,115 divisible by 9
Answer by LinnW(1048)

What is the largest value of X makes this number 577,5X7,115 divisible by 9
If the sum of the digits of a number is divisible by 9 then
the number is divisible by 9
The sum of the digits for 577,5X7,115 = 38 + X
If (38 + X) = 45 then we have one possibility.
38 + X = 45
add -38 to each side
X = 7
Since there is not a single digit positive number greater than 7,
the correct answer must be X = 7
The resulting number is 577,577,115
Notice that 5+7+7+5+7+7+1+1+5 = 45